Both Sue and Joe are students. Your friend Sue has $20,000 of credit card debt (25% interest charge compounded monthly). Sue plans on paying $400 per month over the next 10 years on this credit card. Your other friend Joe has $40,000 of student loan debt (15% interest charge compounded monthly). Joe plans on paying $645.34 each month for the next 10 years.

A) Which person, Sue or Joe, do you feel will pay the most interest expense and why?

B) Which person (Joe or Sue) has the worst debt situation?
Answer #1

a) Sue: PV = 20,000, PMT = -400, rate = 25%/12, N = 120.

use FV function in Excel

amount outstanding in 10 years = 28,698.85

interest paid = 120*400 - 20,000 + 8698.85 = 36,698.85

Joe:

PV = 40,000, PMT = -645.34, rate = 15%/12, N = 120.

use FV function in Excel

amount outstanding in 10 years = 0

interest paid = 645.34 * 120 - 40,000 = 37,440.80

interest paid by Joe is more in the 10 year period but the loan has been settled.

B) Sue is unable to finance her debt and the principal amount has also increased. Her rate of interest is also higher. She is in a worse situation